# Modular memristor model with synaptic-like plasticity and volatile memory

### 05 Mar 2026 v1

@@ Abstract
Compact models of memristors are essential for simulating large-scale neuromorphic systems, yet they often do not include description of complex dynamics like volatile relaxation and synaptic plasticity. We introduce a modular, computationally efficient memristor model that bridges this gap by integrating principles from physics and computational neuroscience. The model defines a framework consisting of a standard formulation of memristive device dynamics, a functional rule mapping state variables to cumulative conductance, a volatility module inspired by the theory of linear viscoelasticity and a saturation module implementing a linear-nonlinear technique. Additionally, we develop a formulation of synaptic-like plasticity inspired by a biological spike-timing-dependent plasticity (STDP) rule, which is compatible with the general framework for memristive devices. Finally, we propose a Laplace transform-based technique to derive the precise form of the mapping from state variables to cumulative conductance, replacing ad hoc voltage-current relationships with principled construction.
@@
We quantitatively validate the complete model against a rich set of experimental data from polymeric memristors exhibiting potentiation, synaptic-like plasticity and volatile decay. Our work presents a new paradigm for memristor modeling that is both practical for large-scale simulation and rich in explanatory power, providing a principled tool for the design of next-generation neuromorphic hardware.
I Introduction
@@
Neuromorphic computing allows physical devices to bring memory and computation closer together, targeting improvements in energy efficiency compared to traditional von Neumann architectures [29, 9]. Memristors, two-terminal devices whose conductance depends on the history of electrical stimulation [2, 1], are particularly well suited to serve as artificial synapses [10]. Their inherent memory, analog tunability, and scalability make them attractive for both fundamental studies and applications in neuromorphic hardware [29].
@@
Compact models such as TEAM and VTEAM [13, 14] provide versatile descriptions of memristive switching [28], widely used in circuit-level simulations. However, for applications in neural networks, additional properties such as volatility and synaptic-like plasticity must be accounted for in a quantitative and flexible manner. The recently proposed V-VTEAM model [19] addresses volatility of conductance by modifying the dynamics of the memristive system, differing from our approach. Volatile conductance relaxation, in particular, enables short-term memory and fading dynamics [27, 11, 26], while spike-timing-dependent plasticity (STDP) underlies learning and long-term adaptation [10, 30, 21].
Biological synapses inspire these features: transient responses decay unless reinforced, whereas repeated pre- and post-synaptic activity leads to long-term modifications. Importantly, STDP can be described in terms of local variables (eligibility traces) that evolve directly from the input spikes, making the mechanism both biophysically plausible and technically practical [6].
To capture memristor volatility, we take inspiration from linear viscoelasticity theory, where material stress depends on the convolution of strain differences with a hereditary kernel [3, 15]. In our case, the internal conductance state depends on the past applied voltage (or equivalently, passed charge) through a convolution with a kernel that decays approximately as $1/t$. This form reflects a broad distribution of relaxation times, consistent with disordered systems and percolation phenomena in which conduction pathways evolve stochastically over time [23, 22].
Our formulation is presented in a voltage-driven form, which is natural for many experimental setups. Nevertheless, thanks to the modular structure of the model, switching to a current-driven variant requires only minor adjustments and does not change the essential dynamics. This flexibility ensures broad applicability across device classes.
We propose a modular memristor model that unifies multiple modules :
@@
    a compact memristive component building on established voltage-based models [24, 14];
@@
    a novel STDP component based on local-variable eligibility traces [6], yielding biologically inspired causal and anti-causal weight updates;
@@
    a cumulative conductance function, establishing the relationship between evolution of state variables and conductance;
@@
    a volatility module inspired by viscoelastic hereditary convolution with a kernel;
@@
    a saturation module, which implements a linear-nonlinear scheme in conjunction with the volatility module to produce the resulting conductance.
A schematic representation of the model is provided in Figure˜1. The model is quantitative, flexible, and computationally efficient. Its modularity allows enabling or disabling volatility and STDP without altering the core equations.
To demonstrate its applicability, we compare model dynamics with experimental data from polymeric memristors developed by Pfleger, Foulger, and co-workers [18, 17, 8, 5]. This device consists of a thin film of poly(N-(3-(9H-carbazol-9-yl)propyl)methacrylamide) sandwiched between ITO and top metallic electrodes. The memristor showed pronounced hysteresis in its I-V characteristics. At higher voltages about +/-5 V it exhibits bistable resistive switching effects, with a maximum ON/OFF ratio of around 200 and good retention time exceeding 2 hours in both ON and OFF states [18]. At lower voltage range the device shows an analog behavior and its resistance, stimulated by a series of low amplitude voltage pulses, can gradually increase or decrease, similarly as the potentiation and depression of neuronal synapses [17].
We have additionally included a technique based on the Laplace transform to determine the functional form of the cumulative conductance function for the specific physical device under evaluation in IV.3. Owing to the simplicity of the model framework, this form could provide insight to the characteristic macroscopic behavior of the device, dependent on the chosen memristor model.
We believe this approach will help bridge the gap between device-level physics and system-level neural computation, with particular relevance for sustainable, polymer-based memristive hardware.
II Model
Our objective is to develop a modular model of a volatile memristor exhibiting spike-timing dependent plasticity (STDP) characteristics. We begin by presenting commonly used voltage-controlled memristor models. Next, we adapt an online synaptic STDP model into a form suitable for non-spiking one-port electronic devices. We draw inspiration from linear viscoelasticity [3] to model the observed volatility in device conductance, introducing a cumulative conductance function together with a decay kernel. Finally, we combine these components and apply a linear-nonlinear technique [16] to both capture the conductance saturation observed in devices and to constrain the admissible conductance range.
A key advantage of our framework is its modularity. One may freely choose a memristor voltage-controlled model, the cumulative conductance function, the decay kernel and saturation function, subject to only minimal constraints, to best match the device being modeled.
@@ II Model
Our approach is inspired by the memristive system framework, but differs from it in several important aspects. Following Eq.˜1, we introduce a vector of four internal state variables: s,w,x,ys,\,w,\,x,\,y, along with a description of their respective dynamics. However, unlike in the formulation above, our conductance function $G$ is not a direct function of the state parameters. Instead, we introduce an intermediate function $H$ and compute the device conductance $G$ using a linear–nonlinear approach combined with the linear viscoelasticity framework.
Namely, we introduce a kernel function $\ker(\,)$ to capture the temporal decay of conductance, a "cumulative conductance" function $H(s,w)$ to describe the non-volatile conductance characteristics and a saturation function $\operatorname{sat}(\,)$ to model conductance saturation. A schematic of the modules’ relationships is provided in Figure˜1.
We note that our model does not require state variable boundedness in the description of the dynamics, in contrast to Strukov et al. [24], Kvatinsky et al. [14]. Instead, we resolve non-physical behavior by the saturation component of the model.
Figure 1: Schematic representation of the model. The model integrates two independent components into the cumulative conductance function $H(s,w)$: voltage-controlled switching dynamics (state variable $s$) providing the memristive core, and STDP-like synaptic plasticity (state variable $w$) implementing learning rules through eligibility traces. The function then undergoes volatile decay through viscoelastic-inspired convolution with a decay kernel $\ker()$. The final conductance $G$ emerges after saturation nonlinearity, reproducing both volatile memory and synaptic plasticity observed in polymeric memristors. This modular approach enables near-independent optimization of each functional component.
II.1 Voltage-controlled model
We present two voltage-controlled memristor models fitting the general framework for memristive devices suitable as the basis for the non-volatile behavior.
A simple and common [24] possibility is the linear voltage controlled memristor model:
@@ II.1 Voltage-controlled model
A more detailed approach due to [14] is also a popular choice:
$\dot{s}=\begin{dcases}k_{\mathrm{off}}\Big(\frac{v(t)}{v_{\mathrm{off}}}-1\Big)^{\alpha_{\mathrm{off}}},&\text{if }0<v_{\mathrm{off}}<v,\\ \makebox[65.46404pt][c]{$0\,,$}&\text{if }v_{\mathrm{on}}<v<v_{\mathrm{off}},\\ k_{\mathrm{on}}\Big(\frac{v(t)}{v_{\mathrm{on}}}-1\Big)^{\alpha_{\mathrm{on}}},&\text{if }v<v_{\mathrm{on}}<0\,.\end{dcases}$
Notably, we have omitted the windowing functions in the above equations, since in our approach the state parameters are unbounded at this stage of the modeling process. Hence at this stage, the so-called linear memristor model of Eq.˜2 is truly a linear differential equation.
@@
II.2 STDP model for a one-port device
In this section, we adapt the local variables synaptic STDP model, as presented in [6], to a one-port physical electronic device, making it compatible with the standard memristive framework. This model provides a phenomenological description of synaptic weight parameter changes based on relative timing of two classes of spiking inputs: presynaptic and postsynaptic. Notably, due to its abstract formulation, the model can be generalized beyond neural systems, provided the two input classes are reinterpreted in an appropriate manner.
@@ II.2 STDP model for a one-port device
Our aim is to reformulate this model for a one-port device, which has important distinctions from biological synapses. The key idea is to replace discrete spike trains with continuous voltage signals, while interpreting one terminal as the ’presynaptic’ terminal and the other as the ’postsynaptic’ terminal. With this identification, the respective spike trains correspond to positive and negative voltages across the device. Accordingly, we substitute the spike trains at each synaptic terminal with the appropriate signed voltage function, as illustrated below:
@@
By treating the variables w,x,yw,\,x,\ y as additional state parameters of the memristor, we can formulate the STDP model for a one-port device as follows, thereby completing the construction.
@@
$\dot{w}$ $=x\,A_{\mathrm{+}}(w)\,v^{-}-y\,A_{\mathrm{-}}(w)\,v^{+}\,.$
We differ from [6] in our approach by considering unbounded $w$ at this stage. As a result the functions $A_{+},\,A_{-}$ do not need to contain a windowing function to force bounds onto the weight parameter.
We note that the exponential decay of the variables $x$ and $y$ is inherited from the synaptic model. However, other volatility dynamics can be considered, for instance by replacing the decay with a more general kernel-based approach. The choice of dynamics determines the shape of the STDP characteristic, in particular the classical double-exponential result is a consequence of the exponential decay for variables $x$ and $y$.
As a final remark, we note that the full memristive dynamics formulation may incorporate coupling by allowing $A_{\mathrm{+}}$ and $A_{\mathrm{-}}$ to depend on $s$.
II.3 Introducing volatility by analogue with linear viscoelasticity
In order to model memristor volatility characteristics, we implement a standard convolution-based approach, inspired by linear viscoelasticity theory [3]. A key advantage is that this approach enables the use of $1/t^{\alpha}$ volatility kernels that are otherwise computationally unavailable. Given the functions $k$ and $H$, denote the Stieltjes convolution as
$k\ast dH\,(t)=\int^{t}_{-\infty}k(t-\tau)\,dH(\tau)$
whenever this Lebesgue-Stieltjes integral is defined [^1] A sufficient condition is that k,Hk,\,H are of bounded variation and vanish for negative times, which is the case for the functions we are considering. In the following, the symbol $\ast$ will always denote the Stieltjes convolution.
In this approach, we need to select a suitable kernel function $\ker()$, which characterizes the device’s conductance volatility, and the cumulative conductance function $H(s,w)$, which describes nonvolatile behavior of conductance. The resulting volatile cumulative conductance $H_{\mathrm{vol}}$ is calculated as the convolution
@@ II.3 Introducing volatility by analogue with linear viscoelasticity
Note that the convolution is calculated with respect to time, since the model state variables $s$ and $w$ are time-dependent. Similarly, the differential $dH$ must also be taken with respect to time, written explicitly [^2] We require of $H$ that it be differentiable almost everywhere.:
@@
This relation motivates the label "cumulative conductance" for the function $H$; it captures the functional relationship of state variables and device conductance in the non-volatile case. Conversely, if $\lim_{t\rightarrow\infty}\ker(t)=0$, we are in the viscoelastic regime, where the system decays to its equilibrium over time.
Our model is therefore sufficiently general to describe devices exhibiting a combination of volatile and non-volatile characteristics.
The form of the kernel and cumulative conductance functions is device-specific and also depends on the chosen voltage-controlled model in Section˜II.1. Inspired by the presentation in Kvatinsky et al. [14], we provide ad-hoc variants for the kernel function and for the cumulative conductance function.
@@
The terminology corresponds to the relationships in the case of the purely elastic regime ($\ker()$ is the Heaviside function), as discussed above. These two versions are generalizations of the current-voltage relationships presented in Kvatinsky et al. [14].
@@
In addition, we offer a rigorous but data-specific approach for deriving the functional forms of the volatility characteristics $\ker()$ and $H$ in Section˜IV.1 and Section˜IV.3, respectively. We do so using data from pulse stimulation of the device, idealized by a single box function. Consequently, in Section˜IV.4 we consider a conductance function of the form
@@
II.4 Introducing saturation via linear-nonlinear approach
@@
II.5 Complete model of volatile memristor with STDP
Combining the results of sections II.1II.4 we formulate the complete model.
  1. 1.
    Consider the vector (s,w,x,y)(s,\,w,\,x,\,y) of state parameters and define the dynamics of the state variables by choosing an equation for $s$ from section II.1, and Eq.˜3 for x,y,wx,\,y,\,w.
  2. 2.
    Next, as per section II.3 choose an appropriate kernel function $\ker()$ and cumulative conductance function $H$ and calculate its derivative.
  3. 3.
    Compute the volatile pre-conductance $H_{\mathrm{vol}}$ by Eq.˜4.
  4. 4.
    The resulting conductance $G(t)$ is obtained from Eq.˜7.
III Experimental data
III.1 Polymeric memristor with neuro-synaptic properties
The experimental validation of our model employs memristive devices based on poly[N-(3-(9H-carbazol-9-yl)propyl)methacrylamide] (PCaPMA). In these devices, a thin PCaPMA film is sandwiched between an indium tin oxide (ITO) bottom electrode and a top electrode made of Au or Al. The resulting structure functions as a two-terminal memristive element whose conductance depends on the history of applied voltage pulses.
IV Fitting model to experimental data
@@
To derive the model for a given device, we first extract the characteristic volatility kernel from the relaxation measurements using an approximation technique. We then apply the Laplace method described in IV.3 to determine the form of the cumulative function and estimate the corresponding parameters. Next, we estimate the parameters of the STDP module from measurements and finally set the saturation module coefficients either to minimally affect the model, or by performing a joint optimization of all parameters.
IV.1 Determining the kernel-function form
@@
IV.2 Finding the kernel-function coefficient $\alpha$
@@
We repeat the above procedure for every measured trace and find an optimal $\alpha$ over all six datasets, as shown in Figure˜3.
Figure 3: Kernel function coefficient estimation. Each trace corresponds to stimulation by pulses over the duration $T_{j}$. For each trace, a fit $g_{j}(t)=b_{j}\,(\frac{1}{(x+\varepsilon)^{\alpha}}-\frac{1}{(x+\varepsilon+T_{j})^{\alpha}})+c_{j}$ is performed using two trace-specific parameters $b_{j}$ and $c_{j}$ together with the global coefficient $\alpha$ to be estimated.
@@
IV.3 Derivation of cumulative conductance
In this section we describe a method for determining the form of the cumulative conductance function $H$. We do so using data from pulse stimulation, again idealized by a single box function similarly as in Section˜IV.2. The procedure yields only the dependence on the parameter $s$, the dependence on $w$ must be copied, or derived in a different way.
Our goal is to find the cumulative conductance function such that the prediction given by the model results in the traces $g_{j}$ from Figure˜3. Note that because the following utilizes the Laplace transform, we must restrict ourselves to positive times, hence in contrast to the discussion in Section˜IV.2 we work with the single box function supported on the interval (0,T)(0,\,T), i.e. $\chi_{(0,\,T)}$.
@@ IV.3 Derivation of cumulative conductance
Figure 4: Cumulative conductance function estimate. Each curve shows the estimated dependence of $\partial_{s}H(s)$ on the state variable $s$, reconstructed from relaxation traces with different stimulation durations $T$ using the procedure in IV.3. The parameters $h_{0}$, $h_{1}$ are chosen so that the traces for all $T$ collapse onto a power-law-like curve with a common a=0.585a=-0.585, well approximated by the parametric form $dH(s)=(h_{1}\,|s|^{a}+h_{0})\,\dot{s}$. This justifies the simple functional form of dHdH used in the main text.
IV.4 Finding cumulative conductance function coefficients
@@
We are considering $dH$ in the form derived in IV.3:
@@
Figure 5: Comparison between the model and the experimental data for a relatively long stimulation input ($T=35$ s). The measured conductance decays after potentiation by a pulse train of total duration $T$ and the solid line shows the corresponding numerical simulation of the model.
Figure 6: Comparison between experimental relaxation and model prediction for a shorter stimulation train ($T=3.5$ s). As in Figure˜2, the dots show the measured post-potentiation conductance decay and the solid line the prediction of the same volatility kernel Eq.˜9 and cumulative conductance function $H(s)$ as in fig.˜5. The overall good agreement across the entire decay window indicates that a single, material-specific kernel captures the volatile dynamics for different stimulation protocols.
IV.5 Modeling STDP
To estimate the coefficients of the STDP component, we extend the previously computed part of the model with the dependence on the weight parameter $w$. Specifically we are now looking to estimate the coefficients from Eq.˜3: $A_{\mathrm{+}},\,A_{\mathrm{-}},\,\tau_{\mathrm{+}},\,\tau_{\mathrm{-}},\,$, as well as the coefficients from the STDP component of $dH$; $h_{01}$ and $h_{00}$.
The data under evaluation consists of three conductance measurements per run. The first determines the steady-state conductance $g_{0}$, while the other two quantify the the conductance change $\Delta G$: one measured 10ms after the end of the second testing pulse (denoted $t_{1}$) and another measured after 100ms (denoted $t_{2}$). We optimize the parameters for the first measurement, since the second one should be a result of the volatility modeled by $\ker()$.
@@ IV.5 Modeling STDP
and estimate the parameters of the model to fit the measurements. The results of the STDP-induced conductance at both sample times are shown in fig.˜7. As a consequence, we verified that the decrease in STDP effects over time is a result of the previously modeled volatility mechanism.
@@
IV.6 Implementing saturation
Up to now, we have ignored the effects of saturation in the modeling for simplicity of presentation. We thus implicitly assumed the effects modeled so far lie in the linear regime of the saturation curve, so that the subsequent addition of the saturation component will have minimal effect.
The above discussed approach may be repeated with chosen coefficients for the saturation module to obtain optimized model parameters that now include the effects of saturation. Suitable coefficients for $g_{\mathrm{max}},\,g_{\mathrm{min}},\,h_{\mathrm{max}}$ and $h_{\mathrm{min}}$ are presented in A.
V Discussion

### 13 Apr 2026 v2

@@ Abstract
Compact models of memristors are essential for simulating large-scale neuromorphic systems, yet they often do not include description of complex dynamics like volatile relaxation and synaptic plasticity. We introduce a modular, computationally efficient memristor model that bridges this gap by integrating principles from physics and computational neuroscience. Starting from standard memristive system dynamics, the framework incorporates synaptic-like plasticity dynamics, a mapping from state variables to cumulative conductance, a volatility module, and a saturation module. The plasticity component is inspired by a biological rule for spike-timing-dependent plasticity (STDP) and is compatible with the general memristive systems formalism. Finally, we propose a Laplace transform-based technique to derive the precise form of the mapping from state variables to cumulative conductance, replacing ad hoc voltage-current relationships with principled construction.
@@
We quantitatively validate the complete model against a comprehensive set of experimental data from polymeric memristors exhibiting potentiation, synaptic-like plasticity and volatile decay. Our work presents a new paradigm for memristor modeling that is both practical for large-scale simulation and rich in explanatory power, providing a principled tool for the design of next-generation neuromorphic hardware.
Teaser: Kernel-based volatility and eligibility-trace plasticity unify short-term memory and learning in polymer memristor compact models.
Introduction
@@
Neuromorphic computing allows physical devices to bring memory and computation closer together, targeting improvements in energy efficiency compared to traditional von Neumann architectures [28, 9]. Memristors are one-port devices whose conductance depends on the history of electrical stimulation [2, 1] and are particularly well suited to serve as artificial synapses [10]. Their inherent memory, analog tunability, and scalability make them attractive for both fundamental studies and applications in neuromorphic hardware [28].
@@
Compact models such as TEAM and VTEAM [13, 14] provide versatile descriptions of memristive switching [27], widely used in circuit-level simulations. However additional properties must be accounted for in a quantitative and flexible manner for applications in neural networks, such as conductance volatility supporting short-term memory and fading dynamics [26, 11, 25] and synaptic-like plasticity enabling learning and long-term adaptation [10, 29, 20]. The recently proposed V-VTEAM model [18] addresses volatility by modifying the dynamics of the memristive system, but a unified framework that combines synaptic-like plasticity, volatility and saturation while remaining modular and computationally efficient is still lacking.
We propose a modular memristor model that augments the standard memristive system description with modules for plasticity, volatility and saturation. This framework unifies:
@@
    a standard memristive systems component utilizing established voltage-driven models [23, 14];
@@
    a novel synaptic-like plasticity component based on local-variable eligibility traces [6], yielding biologically inspired causal and anti-causal weight updates;
@@
    a cumulative conductance function that maps state variables to conductance;
@@
    a volatility module inspired by linear viscoelasticity [3, 15], where material response depends on a hereditary convolution kernel; and
@@
    a saturation module implemented as a linear-nonlinear scheme in conjunction with volatility, to produce bounded device conductance.
One can select the memristive core, cumulative conductance mapping, decay kernel and saturation function to match a given device and enable or disable volatility and/or STDP without changing the underlying modeling framework. Although we present the framework in voltage-driven form consistent with common experimental protocols, its modular structure also permits straightforward translation to current-driven variants.
We validate the approach using experimental data from polymeric memristors developed by Pfleger, Foulger, and co-workers [17, 16, 8, 5]. These memristors consist of a thin film of poly(N-(3-(9H-carbazol-9-yl)propyl)methacrylamide) sandwiched between ITO and top metallic electrodes. The devices studied exhibited pronounced hysteresis in their I-V characteristics, bistable resistive switching effects at higher voltages of +/-5 V (with a maximum ON/OFF ratio of around 200 and retention time exceeding 2 hours)[17], and analog, pulse-driven conductance modulation at lower voltages that resembles potentiation and depression of neuronal synapses [16]. For this class of devices, the internal conductance state depends on the history of applied voltage via a convolution with a kernel that decays approximately as $1/t$. This form reflects a broad distribution of relaxation times, consistent with disordered systems and percolation phenomena in which conduction pathways evolve stochastically over time [22, 21].
Using an appropriate measurement protocol, the distinct effects of the individual components may be effectively separated. To this end, we developed a measurement procedure that exploits the modular structure of the framework to efficiently estimate the model parameters, enabling near-independent optimization of each functional component.
Finally, we introduce a technique based on the Laplace transform to determine the functional form of the cumulative conductance mapping for a given device. Owing to the simplicity of the model framework, this form may provide useful insight to the characteristic macroscopic behavior of the device.
The model reproduces both volatile memory and synaptic plasticity observed in polymeric memristors, while maintaining computational efficiency for large-scale neuromorphic simulations. We believe this approach will help bridge the gap between device-level physics and system-level neural computation, with particular relevance for sustainable, polymer-based memristive hardware.
Results
We present the theoretical model and its application to hardware for parameter fitting.
Model
@@ Results
Our approach is inspired by the memristive system framework, but differs from it in several important aspects. Following Eq.˜1, we introduce a vector of four internal state variables: s,w,x,ys,\,w,\,x,\,y, along with a description of their respective dynamics. To incorporate STDP into our memristor model, we introduce a formulation of spike timing-dependent plasticity dynamics (variables w,xw,\,x and $y$) for one-port electronic devices, compatible with the standard formulation of memristive systems.
However, unlike in the formulation above, our device conductance function $G$ is not directly a function of the state parameters. Instead, we introduce a kernel function $\ker(\,)$ to capture the temporal decay of conductance, a "cumulative conductance" mapping $H(s,w)$ from state variables to non-volatile device conductance and a saturation function $\operatorname{sat}(\,)$ to model device conductance saturation. A schematic of the modules’ relationships is provided in Figure˜1.
Figure 1: Schematic representation of the model. The model combines two independent components into the cumulative conductance $H(s,w)$: voltage-controlled switching dynamics providing the memristive core (state variable $s$) , and STDP-like synaptic plasticity implementing learning rules via eligibility traces (state variable $w$). The cumulative conductance then undergoes volatile decay through viscoelastic-inspired convolution with a decay kernel $\ker()$ and the resulting device conductance $G$ emerges after a saturation nonlinearity.
Memristive system dynamics
This component builds upon established voltage-driven memristive models, suitable as the basis for non-volatile behavior. A simple and common possibility for the memristive model core is the linear voltage controlled memristor model [23]:
@@
A more detailed approach due to [14] is also a popular choice.
Notably, we have omitted the windowing functions in the above equation, because in our approach the state parameters are unbounded at this stage of the modeling process. Hence the so-called linear memristor model of Eq.˜2 is truly a linear differential equation at this stage. We resolve non-physical behavior by the saturation component of the model.
@@
Adaptation of STDP to a memristive device
In this section, we adapt the local variables synaptic STDP model, as presented in [6], to a one-port physical electronic device, making it compatible with the standard memristive systems framework. The model provides a phenomenological description of synaptic weight parameter changes based on relative timing of two classes of spiking inputs: presynaptic and postsynaptic. Notably, due to its abstract formulation, it can be generalized beyond neural systems, provided the two input classes are reinterpreted in an appropriate manner.
@@
Our aim is to reformulate this model for a one-port device, which differs from biological synapses in several important respects. The key idea is to replace discrete spike trains with continuous voltage signals, while interpreting one terminal as the ’presynaptic’ terminal and the other as the ’postsynaptic’ terminal. With this identification, the respective spike trains correspond to positive and negative voltages across the device. Accordingly, we substitute the spike trains at each synaptic terminal with the appropriate signed voltage function, as illustrated below:
@@
Adding variables w,x,yw,\,x,\ y as additional state parameters of the memristive device model, we can formulate the synaptic-like plasticity component for the one-port device as follows:
@@
$\dot{w}$ $=x\,A_{\mathrm{+}}(w)\,v^{-}-y\,A_{\mathrm{-}}(w)\,v^{+}\,$,
where $\tau_{\mathrm{+}},\,\tau_{\mathrm{-}}$ are time constants and the functions $A_{\mathrm{+}},\,A_{\mathrm{-}}$ are weight dependent scaling factors. We denote by $v^{+},\,v^{-}$ the positive and negative parts of $v$, respectively, with the convention $v=v^{+}-v^{-}$.
We differ from [6] in our approach by considering unbounded $w$ at this stage. As a result the functions $A_{+},\,A_{-}$ do not need to contain a windowing function to force bounds onto the weight parameter. We note that in the model, the memristive system components may incorporate coupling by allowing $A_{\mathrm{+}}$ and $A_{\mathrm{-}}$ to depend on $s$.
Volatility kernel and cumulative conductance
In order to model memristor volatility characteristics, we implement a standard convolution-based approach, inspired by linear viscoelasticity theory [3]. A key advantage is that this approach enables the use of $1/t^{\alpha}$ volatility kernels that are otherwise computationally unavailable. We select a suitable kernel function $\ker()$, which characterizes the device’s conductance volatility, and the cumulative conductance function $H(s,w)$, which describes nonvolatile behavior of conductance. The resulting volatile cumulative conductance $H_{\mathrm{vol}}$ is calculated as the convolution
@@
Note that the convolution is calculated with respect to time, since the model state variables $s$ and $w$ are time-dependent. Similarly, the differential $dH$ must also be taken with respect to time, written explicitly:
@@
This relation motivates the label "cumulative conductance" for the function $H$; it captures the functional relationship of state variables and device conductance in the non-volatile case. Conversely, if $\lim_{t\rightarrow\infty}\ker(t)=0$, we are in the viscoelastic regime, where the system decays to its equilibrium over time. Our model is therefore sufficiently general to describe devices exhibiting a combination of volatile and non-volatile characteristics.
The form of the kernel and cumulative conductance functions is device-specific and also depends on the chosen voltage-controlled model. Inspired by the presentation in Kvatinsky et al. [14], we provide ad-hoc variants for the kernel function and for the cumulative conductance function.
@@
The terminology corresponds to the usual relationships in the case of the purely elastic regime ($\ker()$ is the Heaviside function), as discussed above. Thus both versions generalize the current-voltage relationships presented in Kvatinsky et al. [14].
@@
In addition, we offer a rigorous but data-specific approach for deriving the functional forms of the volatility characteristics $\ker()$ and $H$ in Eq.˜9 and Eq.˜11, respectively. We do so using data from pulse stimulation of the device, idealized by a single box function. Consequently, in the fitting section below we consider a conductance function of the form
@@
As a final remark, we note that convolving the kernel with the derivative $dH$ is useful because it allows the convolution term to be effectively ’reset’ by the state parameter derivatives $\dot{s},\,\dot{w}$, which is necessary for the kernel-based approach to work correctly.
Saturation and resulting conductance
@@
Fitting to experimental data
@@
To derive the model for a given device, we first extract the characteristic volatility kernel from the relaxation measurements using an approximation technique. We then apply the Laplace method described below to determine the form of the cumulative function and estimate the corresponding parameters. Next, we estimate the parameters of the STDP module from measurements and finally set the saturation module coefficients either to minimally affect the model, or by performing a joint optimization of all parameters.
Determining the kernel-function form
@@
Finding the kernel-function coefficient $\alpha$
@@
We repeat the above procedure for every measured trace and find an optimal $\alpha$ over all five datasets, as shown in Figure˜3.
Figure 3: Kernel function coefficient estimation. Each trace corresponds to stimulation by pulses over the duration $T_{j}$. For each trace, a fit $g_{j}(t)=b_{j}\,(\frac{1}{(t+\varepsilon)^{\alpha}}-\frac{1}{(t+\varepsilon+T_{j})^{\alpha}})+c_{j}$ is performed using two trace-specific parameters $b_{j}$ and $c_{j}$ together with the global coefficient $\alpha$ to be estimated.
@@
Derivation of cumulative conductance
In this section we describe a method for determining the form of the cumulative conductance function $H$. We do so using data from pulse stimulation, again idealized by a single box function similarly as above. The procedure yields only the dependence on the parameter $s$, the dependence on $w$ must be copied, or derived in a different way.
Our goal is to find the cumulative conductance function such that the prediction given by the model results in the traces $g_{j}$ from Figure˜3. Note that because the following utilizes the Laplace transform, we must restrict ourselves to positive times, hence in contrast to the discussion above we work with the single box function supported on the interval (0,T)(0,\,T), i.e. $\chi_{(0,\,T)}$.
@@
Figure 4: Cumulative conductance function estimate. Each curve shows the estimated dependence of $\partial_{s}H(s)$ on the state variable $s$, reconstructed from relaxation traces with different stimulation durations $T$ using the discussed procedure. The parameters $h_{0}$, $h_{1}$ are chosen so that the traces for all $T$ collapse onto a power-law-like curve with a common a=0.585a=-0.585, well approximated by the parametric form $dH(s)=(h_{1}\,|s|^{a}+h_{0})\,\dot{s}$. This justifies the simple functional form of dHdH used in the main text.
Finding cumulative conductance function coefficients
@@
We are considering $dH$ in the form derived in Eq.˜11:
@@
Figure 5: Comparison between the model and the experimental data for stimulation input of $T=35$ s and $T=3.5$ s. The measured conductance decays after potentiation by a pulse train of total duration $T$ and the solid line shows the corresponding numerical simulation of the model. The overall good agreement across the entire decay window indicates that a single, material-specific kernel captures the volatile dynamics for different stimulation protocols.
Modeling STDP
To estimate the coefficients of the STDP component, we extend the previously computed part of the model with the dependence on the weight parameter $w$. Specifically we are now looking to estimate the coefficients from Eq.˜3: $A_{\mathrm{+}},\,A_{\mathrm{-}},\,\tau_{\mathrm{+}},\,\tau_{\mathrm{-}}\,$, as well as the coefficients from the STDP component of $dH$; $h_{01}$ and $h_{00}$.
The data under evaluation consists of three conductance measurements per run. The first determines the steady-state conductance $g_{0}$, while the other two quantify the conductance change $\Delta G$: one measured 10ms after the end of the second testing pulse (denoted $t_{1}$) and another measured after 100ms (denoted $t_{2}$). We optimize the parameters for the first measurement, since the second one should be a result of the volatility modeled by $\ker()$.
@@
and estimate the parameters of the model to fit the measurements. The results of the STDP-induced conductance at both sample times are shown in fig.˜6. As a consequence, we verified that the decrease in STDP effects over time is a result of the previously modeled volatility mechanism.
@@
Implementing saturation
Up to this point, we have ignored the effects of saturation in the modeling for simplicity of presentation. We thus implicitly assumed the effects modeled so far lie in the linear regime of the saturation curve, so that the subsequent addition of the saturation component will have minimal effect.
The above discussed approach may be repeated with chosen coefficients for the saturation module to obtain optimized model parameters that now include the effects of saturation. Suitable coefficients for $g_{\mathrm{max}},\,g_{\mathrm{min}},\,h_{\mathrm{max}}$ and $h_{\mathrm{min}}$ are presented in the Appendix.
Methods
Experimental data
Polymeric memristor with neuro-synaptic properties
The experimental validation of our model employs memristive devices based on poly[N-(3-(9H-carbazol-9-yl)propyl)methacrylamide] (PCaPMA). In these devices, a thin PCaPMA film is sandwiched between an indium tin oxide (ITO) bottom electrode and a top electrode made of Au or Al. The resulting structure functions as a one-port memristive element whose conductance depends on the history of applied voltage pulses.
Discussion